### v1

# Surface-mediated reduction of radiation damage in tungsten revealed by advanced ion channeling analysis

Surface-mediated reduction of radiation damage in tungsten revealed by advanced ion channeling analysis
Tungsten is a leading candidate material for plasma-facing components in future fusion reactors. Extensive studies have been performed to better understand its behavior under irradiation. Recent experiments of Rutherford backscattering spectrometry in channeling mode indicated a marked reduction in radiation damage in single-crystal tungsten samples irradiated by self-ions when the irradiation temperature was increased from room temperature to 800 K. However, the underlying mechanism for this damage reduction remains unclear. In this work, by combining Rutherford backscattering spectrometry in channeling mode and molecular dynamics simulations, we identify a pronounced surface effect at elevated temperatures, characterized by a significant reduction of dislocation density near the surface. We demonstrate how our unique analysis method can clearly resolve a dislocation-free zone and a transition region with suppressed defect density before reaching the bulk value. The strong surface effect at elevated temperatures is explained by considering the coherent drift motion of dislocation loops towards the surface, highlighting alternative perspectives on mitigating radiation damage by increasing dislocation mobility.
The extreme operational conditions of future fusion reactors demand plasma-facing materials that can withstand intense heat fluxes and severe radiation damage [Bolt2002, Schmid2024]. Currently, one of the leading candidates for plasma-facing materials is tungsten due to its highest melting point of all metals, low sputtering yield, and low tritium retention [Philipps2011, Rieth2013]. Under irradiation, high-energy particles can displace tungsten atoms, causing damage that alters the properties of the material. Hence, extensive research has been conducted on radiation effects in tungsten, with particular emphasis on defect production and evolution [Yi2016, Atwani2018, Mason2021, Wang2023a, Markelj2024].
To understand these irradiation-induced phenomena, a thorough characterization of irradiation-induced defects is essential, but remains non-trivial. Recently, a comprehensive experimental study of self-ion irradiation of single-crystal tungsten showed that, at a damage level of 0.02 dpa, increasing the irradiation temperature from 290 K to 800 K significantly lowers the signals of Rutherford backscattering spectrometry in channeling mode (RBS/c) [Zavasnik2025]. RBS/c is a special ion beam analysis technique that is very sensitive to several types of crystal defects [Gemmell1974, Feldman1982, Zhang2015c]. Therefore, the pronounced decrease in the RBS/c signal indicates a substantial reduction in radiation damage at 800 K. Understanding the mechanism underlying this reduction in radiation damage is essential, since future fusion reactors will operate at elevated temperatures [Bolt2002]. Nevertheless, it is difficult to unambiguously determine the nature of the radiation defects formed in tungsten solely on the basis of experiments without advanced simulations.
In this Letter, we reveal the mechanisms behind the reduction of radiation damage in tungsten, as observed in the RBS/c experiments [Zavasnik2025], using state-of-the-art simulation methods. We characterize the depth distribution of radiation defects, particularly dislocation loops, and demonstrate how a surface can generate a pronounced effect on the dislocation distribution at elevated temperatures. We demonstrate that this strong surface effect can be explained by temperature enhanced drift motion of dislocation loops, offering new insights into the relationship between defect mobility and resistance of materials to radiation damage.
A series of RBS/c experiments [Zavasnik2025] were performed on the -oriented single-crystal tungsten samples irradiated with 10.8 MeV tungsten ions at different conditions (for details of the experiments, see Section 1 of Supplementary Materials). In this work, we chose to interpret the specific RBS/c spectra generated from the tungsten samples irradiated to 0.02 dpa at 290 K and 800 K, in which the energy of probing ions, i.e., 4He ions, was 3 MeV. For this purpose, we performed RBS/c simulations using the RBSADEC code [Zhang2016, Jin2020, RBSADEC_ND], which can generate signals directly comparable to experiments. A critical prerequisite in the simulation is to supply the RBSADEC code with simulated targets accurately representing realistic atomic structures. To this regard, molecular dynamics (MD) simulations were performed to mimic radiation effects in the experiments. Radiation damage was created by overlapping collision cascades initiated by a 10 keV primary knock-on atom, similarly to Ref. Granberg2021. After the generation of radiation damage, dozens of MD cells were connected along the direction and were relaxed in an additional MD simulation, which finally produced a single super-cell, whose size along the direction is 1.4 m. Our previous study [Markelj2024] demonstrated that simulated RBS/c spectra from this type of super-cell can well reproduce experimental signals at 0.02 dpa and at room temperature (RT). In this study, we further improved the connection process leading to that all MD cells can be perfectly connected.
When connecting the MD cells, each MD cell can have different numbers of collision cascades and hence different levels of radiation damage. The depth profile of cascade number can be determined in such a way that the depth profile of damage dose in dpa corresponds to that calculated by the SRIM software, which is based on the binary collision approximation (BCA) [Ziegler_ND, Ziegler2015]. Hereafter, we refer to this methodology as the BCA-MD approach for the super-cell generation. An alternative approach involves manually adjusting the depth profiles of MD cells to directly fit the experimental spectra. Hereafter, we refer to this approach as the Fit-MD approach (For details of RBS/c and MD simulations, see Section 1 and Section 2 of Supplementary Materials).
Fig. 1(a) shows the experimental and simulated RBS/c signals obtained by using 3 MeV He ion beams on -oriented tungsten samples, which are represented by the dots and lines, respectively. The simulations are based on the BCA-MD approach. The simulated result at RT agrees very well with the experimental one, with only a small yield overestimation in the high energy region (corresponding to the surface region) as shown in the inset. The yield of the experimental RBS/c spectrum of the sample irradiated at 800 K is almost half that of the sample irradiated at RT, as shown in Fig. 1(a). This substantial decrease in RBS/c signals provides a strong indication that irradiation at elevated temperatures results in a reduction in radiation damage. Furthermore, at this elevated temperature, we observe a significant discrepancy between the simulated and experimental spectra. In fact, the simulated spectra at 800 K remain almost the same as those at RT. To understand the difference in the radiation effects in tungsten irradiated at different temperatures, we quantitatively analyzed the atomistic structures obtained by means of BCA-MD with predominant radiation defects, namely 1/2 dislocation loops, which are commonly reported to form in tungsten under ion irradiation [Ferroni2015, Durrschnabel2021, Markelj2024].
Fig. 1(b) shows that the dislocation density at RT practically remains constant ( m-2) throughout the damaged layer up to 860 nm beneath the surface, then gradually decreases to zero by 1.3 m. (The depth resolution in this study is determined by the size of individual MD cells, i.e., 20 nm). For the density distribution at 800 K, unlike the significant decrease of RBS/c signals, the general trend at 800 K is similar to that at RT, except for the slightly lower dislocation density within the damaged region ( m-2). Hence, the discrepancy between the simulated and experimental RBS/c spectra measured at RT and at 800 K is likely to stem from a factor not considered in the BCA-MD approach, for example, significant defect mobility at long time scales, inaccessible with conventional MD.
Using the Fit-MD approach, the simulated signals can perfectly reproduce the experimental ones at both temperatures, as shown in Fig. 2(a). In addition, we also fitted the pristine spectrum; for more details, see Section 2 of Supplementary Materials. The depth profiles of the dislocation density are presented in Fig. 2(b). In these results, we observe that the maximum of the dislocation density in the damaged samples, see Fig. 2(b), remains approximately the same as in the BCA-MD simulation cells shown in Fig. 1(b). However, the variations in depth distributions of the dislocation density are much stronger in the Fit-MD approach. Note that the fitting of RBS/c spectra depends on two parameters: dislocation density and loop size. Nonetheless, we emphasize that the shape of density profile is the most critical factor for our subsequent analysis. For details of dislocation size calculations, see Section 3 of Supplementary Materials.
Figs. 3(a) and 3(b) compare the depth distribution of the dislocation density, which we obtained from the two approaches at RT and 800 K, respectively. The Fit-MD results presented in Fig. 3 are obtained by subtracting the density in the pristine sample from the damaged ones given in Fig. 2(b), thus excluding the influence of the signal from the pristine samples. At RT, both approaches yielded similar results with extended plateaus in the central regions of the depth profile. However, the Fit-MD results indicate that the defect density must decrease near the surface (< 200 nm). Indeed, the formation of a dislocation-denuded zone of 10 nm was observed in the TEM studies at RT [Zavasnik2025]. This is consistent with calculations based on elasticity theory [Jager1975] showing that when the distance between dislocation loops and the free surface is only a few times larger than the loop size, the surface is effective in attracting and annihilating dislocation loops.
Fig. 3 (b) shows that, at 800 K, there are significant differences between the two approaches. Although the result of the BCA-MD approach still shows a nearly constant dislocation density that extends from the surface to the bulk region ( 900 nm), the result of the Fit-MD approach indicates a pronounced surface effect. We observe a dislocation-free zone with a thickness of 83 nm, which is much larger than the size of the dislocation loops. Conventionally, surface effects are mainly characterized by the width of such a defect-free zone. However, since the defect-free zone must be followed by a region where the defect density progressively increases from zero to the bulk value, the reduction in the defect density in this transition region should also be attributed to the surface effect [Vigeholm1965, LiYong2023]. In our case, this transition region covers an extended depth, ranging from 83 nm to at least 370 nm, in which the dislocation density is substantially lower than that in the bulk region, where the defect distribution is nearly constant. A key challenge in characterization of the transition region lies in determination of the onset of the bulk region. By comparing the defect distribution in the two approaches, we are able to delineate the clear boundary between the transition and bulk regions ( 370 nm), thus highlighting the capability of our RBS/c analysis. In the previous TEM study [Zavasnik2025], the depth distribution of the total dislocation density (including both lines and loops) did not clearly reveal such a strong surface influence. However, if we focus on the contribution given solely by dislocation loops, we can find that their distribution is also significantly affected by the surface at 800 K, showing a similar trend as observed in our RBS/c results.
By extending to nearly a third of the damaged region (370 nm) that was produced in tungsten by the irradiation of 10.8 MeV self-ions, the surface effect appears to play a crucial role in determination of radiation defects in tungsten at 800 K. Even the thickness of the dislocation-free zone (83 nm) is not a negligible value. This dislocation-free zone is much larger than the denuded zone observed by TEM at RT (10 nm) [Zavasnik2025], and is comparable or even greater than the grain size (35 - 100 nm) in a nanocrystalline tungsten samples, which show good irradiation tolerance [Atwani2018, Atwani2019a]. To better understand the mechanisms underlying this pronounced surface effect, in particular, the formation of a dislocation-free zone, we conducted calculations to quantify how the surface affects dislocation loops. Fig. 4(a) presents the stress on a 1/2 circular dislocation loop in tungsten as a function of the distance between the loop and the open surface. The loop of 5 nm in diameter is parallel to the surface and its Burgers vector is parallel to the surface normal. The stress is calculated based on Baštecká’s equation [Bastecka1964, Fikar2015]. Conventionally, a defect-free zone is determined by comparing the stress applied to the loop with a critical slip stress [Bastecka1964], such as the critical resolved shear stress (CRSS) [Jager1975a]. By comparing the stress exerted at the loop in Fig. 4(a) with CRSS of tungsten obtained from experiments [Brunner2000], we estimate the dislocation-free zone to be 5 and 13 nm at RT and 800 K, respectively. Although the former is close to the TEM observation (10 nm), the latter is much lower than that obtained from our RBS/c analysis (83 nm). In fact, when the distance between the loop and the open surface is 83 nm, the stress on the loop ( 0.01 MPa) is very small compared to the CRSS at 800 K (13 MPa). Most likely, this discrepancy arises from the fact that the CRSS in BCC metals is mainly determined by screw dislocations [Ohsawa2005, Clouet2021]. However, the irradiation-induced dislocation loops are mainly of edge character, and studies indicate that the mobility of edge dislocations is higher than that of screw dislocations [Solomon1971, Vitek1973].
Hence, instead of relying on a single critical stress to evaluate the dislocation-free zone, we propose to focus on the dynamic behavior of dislocation loops under the influence of external stress, i.e., drift motion of dislocation loops [Hirth1982, Mehrer2007]. Under the force induced by the surface effect, , the drift velocity of dislocation loops, , can be calculated as follows:
where is the loop mobility, which depends on the temperature, , the Burgers vector of the loop, , the loop diameter, , the external stress, , and the effective migration energy, . Note that the drift velocity increases as the loop approaches the open surface, due to the higher surface force. Using eV (from dislocation annealing experiments [Ferroni2015]), we can calculate that a loop takes 63.6 s to cross the dislocation-free zone (83 nm) at 800 K. For calculation details, see Section 4 of Supplementary Materials. In ion-irradiation experiments, reaching a damage dose of 0.02 dpa typically requires an irradiation time on the order of 100 s [Markelj2024]. The similarity between these timescales supports that the formation of large dislocation-free zone can be explained by the drift motion of loops at high temperature.
Fig. 4(b) demonstrates the rapid decrease of the time needed for a dislocation loop to cross the dislocation-free zone (83 nm) with a higher irradiation temperature. This time also decreases with a smaller effective migration energy. The effective migration energy in Eq. (1) should represent an average value. Therefore, dislocation loops with smaller migration energies can traverse a greater distance before reaching the surface, contributing to the formation of the transition region. Experimental studies [Yao2008, Yi2015a] on BCC metals indicate that the surface effect should also depend on the surface orientation, with the surface being the most efficient at absorbing 1/2 loops. Since our sample surface is also , this orientation further amplifies the surface effect.
Defect mobility has been recognized as a key parameter that can be tuned by specific material design to improve radiation resistance [Granberg2016, Atwani2019, Nori2025]. Although previous studies have shown that suppressing dislocation mobility reduces radiation damage in highly concentrated alloys, such as multicomponent concentrated single phase alloys [Granberg2016], our work demonstrates that increased dislocation mobility drives the thermally enhanced surface effect observed here. Thus, the preference for higher or lower defect mobility in achieving radiation tolerance is expected to depend critically on the intrinsic properties of the material and the prevailing operating conditions.
In conclusion, we investigated the reduction of experimental RBS/c signals in single-crystal tungsten irradiated at elevated temperatures by combining advanced RBS/c and MD simulation methods. Our analysis clearly shows that raising the irradiation temperature from room temperature to 800 K significantly lowers the near-surface dislocation density. This pronounced defect reduction can be attributed to a thermally enhanced surface effect, driven by the higher mobility of dislocation loops at elevated temperatures, which promotes the drift motion of loops towards the surface.
This work was partially carried out within the framework of the EUROfusion Consortium, funded by the European Union via the Euratom Research and Training Programme (Grant Agreement No 101052200 — EUROfusion). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the European Commission can be held responsible for them. The authors wish to acknowledge CSC – IT Center for Science, Finland, for computational resources, and the Finnish Computing Competence Infrastructure (FCCI) for supporting this project with computational and data storage resources. Xin Jin thanks to Shenzhen Science and Technology Program (ZDSYS20230626091501002).